OpenStudy (anonymous):

forgot how to simplify equations like 40((1/x)-(1/(x+40)))= 1/3 and (1000/x)-(1000/(x-120)) = 1/2 Please help!

OpenStudy (saifoo.khan):

WOW, thts \[\Huge HUGE\] lol.

OpenStudy (anonymous):

you're cute :-) I forgot how to simplify such equations too :-(

OpenStudy (saifoo.khan):

Lol, i thought u knew the answer.. well, asker.. please write the ques more clearly.. i will solve it for u.. or join me in my class.

OpenStudy (anonymous):

plugging in those equations in wolf gives really tough answers :-P

OpenStudy (anonymous):

someone....? please give me an answer? I'm really stuck

OpenStudy (anonymous):

errrm... yeah I think I get it but...still a bit confused... don't worry then

OpenStudy (anonymous):

Trust me, everybody here is as confused as you are about solving such equations.

OpenStudy (anonymous):

40((1/x)-(1/(x+40)))= 1/3 40/x -40/(x+40) = 1/3 (40(x+40)-40x) / x(x+40) = 1/3 1600/x(x+40) = 1/3 4800= x^2 +40x x^2+40x-4800 = 0 quadratic.